A student has a 10 cm3 sample of 1.00 × 10–2 mol dm–3 methanoic acid solution. The student is asked to dilute the methanoic acid solution to a concentration of
2.00 × 10–4 mol dm–3 by adding distilled water.
Which volume of water should be added?
1
Expert's answer
2020-07-07T14:18:45-0400

10 mL = 0.01 L

0.01 x 1.00 × 10–2 = 1.00 × 10–4 mol of solute in the stock solution

C = n / V

New volume of the solution would be: V = n / C

V = 1.00 × 10–4 / 2.00 × 10–4

V = 0.5 L

0.5 L - 0.01 L = 0.49 L of water to add.